Managing Short-Term Resources and Obligations Situation 2. Saint Joseph Hospital buys 600 boxes of latex gloves every year. Each box costs the hospital $8. The cost to place an order for the gloves-which covers the employee staff time, shipping costs, the hospitals receiving center for inventorying, for example-is estimated at S30 per order Carrying costs includes out-of-pocket costs and capital costs. Out-of-pocket costs related to carry costs (for storing the boxes, verifying the inventory periodically, etc.) are estimated at S2 per box per year. To calculate capital costs per box related to carrying costs, the hospital use a 3 percent interest rate Question 1. Calculate capital costs per box (Show how you calculate it) Question 2. Calculate a total carrying cost per box per year (Show how you calculate it) Question 3. How many boxes should be ordered at a time (Show how you calculate it)?
Answer #1

ANSWER - CAPITAL COST = One-time setup cost of a plant or project, after which there will only be recurring operational or running costs.
SO, TOTAL CAPITAL COST = $ 8 *600 = $ 4800 + OUT OF POCKET COST ($2 * 600= $$1200) = $6000

COST PER BOX = $6000/ 600 = $ 10 PER BOX

CAPITAL COST = $ 10 + $10 *0.03 (INTEREST COST) = $ 10.3 PER BOX

NO OF ORDERS = 600 /30 = 20 ORDERS PER YEAR

AND TOTAL COST OF ORDERS = 20 ORDERS *COST PER ORDER $ 30 = $ 600.

THIS IS THE WAYT TO CALCULATE THE Q1, 2 AND 3.